OpenStudy (anonymous):

Write the equation of the circle with center (3, 2) and radius r = 7. Use the ^ key for the exponents. Write your answer as the example: (x - 4)^2+(y+8)^2=25

OpenStudy (dumbcow):

(x-3)^2 + (y-2)^2 = 49

OpenStudy (anonymous):

in general, equation of a circle with center at A,B and radius R is (x-A)^2+(y-B)^2 = R^2